OpenStudy (anonymous):

solve !!! xsquared=144

OpenStudy (bibby):

What is the inverse of squaring a number? Taking the square root of that thing. Take the square root of both sides to get x = sqrt(144)

OpenStudy (anonymous):

i got 12

OpenStudy (bibby):

That's the right answer. You can check by plugging back in. 12*12=144

OpenStudy (anonymous):

thanks can we do another?

OpenStudy (bibby):

Sure

OpenStudy (anonymous):

what about something like Xsquared -8=0

OpenStudy (bibby):

Like other equations, you want to isolate the x before anything. So you'd add 8 to both sides. Then you get x^2=8 take the square root of both sides and you get x = sqrt(8) or 2sqrt(2)

OpenStudy (anonymous):

i got 2 is that right?

OpenStudy (bibby):

well 2squared is 4 so no, taking the square root of x^2 gives you x. Taking the square root of 8 gives you \[ \sqrt{4} \sqrt{2} aka ... 2\sqrt{2}\]